40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 2 01:18:30 2014, Jenne, Update, LSC, Attempt at PRMI+2 arms 
    Reply  Thu Apr 3 00:30:52 2014, Jenne, Update, LSC, Attempt at PRMI+2 arms 
       Reply  Thu Apr 3 03:06:34 2014, Jenne, Update, LSC, Attempt at PRMI+2 arms 
          Reply  Thu Apr 3 18:09:54 2014, ericq, Update, LSC, Attempt at PRMI+2 arms offsetTRY.pdf2dhist.pdfpdh_trans.pdf
Message ID: 9774     Entry time: Wed Apr 2 01:18:30 2014     Reply to this: 9780
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Attempt at PRMI+2 arms 

I was trying to lock PRMI+2 arms, but I'm losing patience with the MC losing lock.  It's mostly well-behaved, but every now and again it'll lose lock, and it always seems to be just when I've gotten to a delicate part.  Anyhow.  I don't think anything needs fixing.

I am not able today to acquire PRMI lock with REFL 165 I&Q, nor am I able to follow Koji's prescription to transition to 165 from 55.  However, I am able to transition to, or just straight-up aquire on, REFL 33 I&Q.  The new ASC is awesome.

Before trying to lock with REFL165, I redid it's demod phase, by actuating on the PRM while locked on REFL 55, and minimizing the PRCL signal in the Q-phase.  Old 165 phase was -83.5 degrees, new is -79.5 degrees.

I then tried the procedure of misaligning the PRM, acquiring ALS lock of both arms, finding the arm IR resonances, and moving the arms off resonance.  Then I restored the PRM, and locked the PRMI on sidebands using REFL33 I&Q. 

Something was a little weird and unexpected:  If the arms were not far, far off resonance, there was a large MICH offset, so that AS was bright, and POP was pretty dark.  If I moved the arms farther from resonance, POP would come back to the "normal" brightness, which was ~460 counts on POP110I.  When POP was dark-ish, POP110I was about 150, although this number could be changed by moving the arms around.  This number was not dependent on PRM alignment.  Also, PRCL was not yet starting to resonate the carrier, because POPDC stayed at its low sideband-lock level of about 90 counts (vs. 2,000+ for a PRMI-only carrier lock), and POP was visually dark on the camera. 

---------------------------------

While playing with PRMI + 2 ALS arms is entertaining for the evening, I don't yet have a plan for dealing with the changing CARM optical plant for IR signals situation.  That's in progress in the daytime.

ELOG V3.1.3-